You have been given the following information about the production of Usher Co., and are asked to provide the plant manager with information for a meeting with the vice president of operations.

Standard Cost Card
Direct materials (6 pounds at $5 per pound) $30.00
Direct labor (0.70 hours at $10) 7.00
Variable overhead (0.70 hours at $4 per hour) 2.80
Fixed overhead (0.70 hours at $9 per hour) 6.30
$46.10


The following is a variance report for the most recent period of operations.

Variances

Costs

Total Standard Cost

Price

Quantity

Direct materials $405,000 $4,141 F $9,100 U
Direct labor 94,500 5,599 U 7,300 U

How many units were produced during the period?

Number of units

eTextbook and Media

How many pounds of raw materials were purchased and used during the period?

Raw material pounds

eTextbook and Media

What was the actual cost per pound of raw materials? (Round answer to 2 decimal places, e.g. 1.25.)

Actual cost per pound of raw materials

$

/lb

eTextbook and Media

How many actual direct labor hours were worked during the period?

Actual direct labor hours hours

eTextbook and Media

What was the actual rate paid per direct labor hour? (Round answer to 2 decimal places, e.g. 1.25.)

Actual rate paid per direct labor hour

$

/hr
Answer #1

Answer a)

Calculation of number of units produced during the period

Standard Cost of Direct material = (Standard quantity of raw material for actual output) X (Standard price per pound of direct material)

Where,

Standard quantity of raw material for actual output = (Standard quantity of raw material per unit of output) X (Actual number of units produced)

Putting the value of Standard quantity of raw material for actual output in standard cost of direct material formula we get:

Standard Cost of Direct material = (Standard quantity of raw material per unit of output X Actual number of units produced) X (Standard price per pound of direct material)

$ 405,000 = (6 pounds X Actual number of units produced) X $ 5.00 per pound

405,000/5 = (6 X Actual number of units produced)

81,000 = (6 X Actual number of units produced)

Actual number of units produced = 81,000/ 6

                                                            = 13,500 units

Therefore the company actually produced 13,500 units during the period.

Summary of Answers for the following parts:

Particulars

Standard

Actual

Quantity

Unit Cost

Total Cost

Quantity

Unit Cost

Total Cost

Direct Material

81,000 pounds

$ 5.00

$ 405,000

82,820 pounds

$ 4.95

$ 409,959

Direct Labour

9,450 hours

$ 10.00

$ 94,500

10,180 hours

$ 10.55

$ 107,399

Answer b)

Calculation of quantity of raw material purchased and used during the period

Material usage variance = (Standard Quantity – Actual Quantity) X Standard price per unit of raw material

- $ 9,100 = (81,000 pounds – Actual quantity of raw material) X $ 5.00 per pound

- 9,100/ 5.00 = (81,000– Actual quantity of raw material)

-1,820 = (81,000 – Actual quantity of raw material)

Actual quantity of raw material = 81,000 + 1,820

                                                    = 82,820 pounds

Therefore the actual quantity of raw material purchased and used is 82,820 pounds.

Working notes:

Standard quantity of raw material for actual output:

Standard quantity of raw material for actual output = Total units produced X Standard raw material quantity per unit of output

                                                                 = 13,500 units X 6 pounds per unit

                                                                = 81,000 pounds

Therefore the standard quantity of raw material for actual output of 13,500 units is 81,000 pounds           

Answer c)

Calculation of actual cost per pound of raw material

Material Price variance = (Standard Price per unit of raw material – Actual cost per unit of raw material) X Actual Quantity of raw material used

$ 4,141 = ($ 5.00 - Actual cost per unit of raw material) X 82,820 pounds

4,141/82,820 = (5.00 - Actual cost per unit of raw material)

0.05 = (5.00 - Actual cost per unit of raw material)

Actual cost per unit of raw material = (5.00 – 0.05)

                                                                = $ 4.95 per pound

Therefore the actual cost per pound of raw material is $ 4.95.

Answer d)

Calculation of actual direct labour hours worked during the period

Labour efficiency variance = (Standard direct labour hours – Actual direct labour hours) X Standard rate per direct labour hour

- $ 7,300 = (9,450 direct labour hours – Actual direct labour hours) X $ 10.00 per direct labour hour

- 7,300/ 10.00 = (9,450– Actual direct labour hours)

-730 = (9,450 – Actual direct labour hours)

Actual direct labour hours = 9,450 + 730

                                                = 10,180 labour hours

Therefore the actual direct labours hours are 10,180.

Working notes:

Standard direct labour hours for actual output:

Standard direct labour hours for actual output = Total units produced X Standard direct labour hours per unit of output

                                                                 = 13,500 units X 0.70 hours per unit

                                                                = 9,450 hours

Therefore standard direct labour hours for actual output of 13,500 units are 9,450 hours.

Answer e)

Calculation of actual direct labour cost paid per direct labour hour

Labour rate variance = (Standard rate per direct labour hour – Actual rate per direct labour hour) X Actual labour hours

-$ 5,599 = ($ 10.00 - Actual rate per direct labour hour) X 10,180 labour hours

- 5,599/10,180 = (10.00 - Actual rate per direct labour hour)

-0.55 = (10.00 - Actual rate per direct labour hour)

Actual rate per direct labour hour = 10.00 + 0.55

                                                                = $ 10.55 per direct labour hour

Therefore the actual direct labour cost paid is $ 10.55 per direct labour hour